OpenStudy (anonymous):

A turtle and a rabbit are in a race to see who is first to reach a point 100 feet away. The turtle travels at a constant speed of 20 feet per minute for the entire 100 feet. The rabbit travels at a constant speed of 40 feet per minute for the first 50 feet, stops for 3 minutes, and then continues at a constant speed of 40 feet per minute for the last 50 feet. Determine which animal won the race and by how much time.

OpenStudy (ujjwal):

can you calculate the time taken by turtle to cover the whole distance of 100 m?

OpenStudy (anonymous):

i dont know what you mean...

OpenStudy (ujjwal):

Can you give formula for speed?

OpenStudy (anonymous):

as in do i know it or can i use it

OpenStudy (ujjwal):

both

OpenStudy (anonymous):

no i dont know and i can use it as long as i get the right answer !

OpenStudy (callisto):

For speed, it's a rate... the formula is \[speed = \frac{distance \ travelled }{time}\] Got it so far?

OpenStudy (anonymous):

im confused about the rabbit stopping for 3 minutes

OpenStudy (callisto):

Can we do it step by step?

OpenStudy (anonymous):

sure!

OpenStudy (callisto):

First, do you understand that formula?

OpenStudy (anonymous):

yes

OpenStudy (callisto):

Now, consider turtle first, ''constant speed of 20 feet per minute for the entire 100 feet.'' So, speed = 20 ft/min and the distance travelled = 100 feet By applying \[speed = \frac{distance \ travelled}{time}\]Can you calculate the time turtle used to finish the race?

OpenStudy (anonymous):

5 mins?

OpenStudy (callisto):

Yes Now, consider rabbit~ There are 3 stages Stage 1: constant speed of 40 feet per minute for the first 50 feet Stage 2: stops for 3 minutes Stage 3: constant speed of 40 feet per minute for the last 50 feet. Consider stage 1, How much time did it spend?

OpenStudy (anonymous):

1.25

OpenStudy (anonymous):

do you add 1.25+1.25+3

OpenStudy (callisto):

Yes, for the time taken of rabbit in the race

OpenStudy (anonymous):

so the rabbit took 5.50 mins and the turtle took 5 mins. Is the answer the turtle and by 50 seconds ?

OpenStudy (anonymous):

.50 *

OpenStudy (callisto):

Not really... 5.5 minutes = 5 minutes 30 seconds

OpenStudy (anonymous):

oops ! so the answer is the turtle by 30 seconds

OpenStudy (callisto):

lol It should be turtle won the race by 30 seconds

OpenStudy (anonymous):

okiedokie thank you !! you're awesome! =)

OpenStudy (callisto):

welcome :)

OpenStudy (callisto):

Determine which animal won the race and by how much time. I suppose you need to write the difference of the turtle's time and rabbit's time...

OpenStudy (anonymous):

sorry that didnt even make sense what i wrote ! how did you know that .50 is 30 seconds

OpenStudy (callisto):

1 minute = 60 seconds. 0.5 minute = 0.5 (60 seconds) = 30 seconds

OpenStudy (anonymous):

wow im stupid ! thanks again !

OpenStudy (callisto):

Welcome~ And you're not stupid!!!!